solve the equation: 6(W + 7) - 3 (W + 7) = 0 , my answer is -42, is this right
if not please show me step by step how to do this problem
thnk you

Answer by Alan3354(69443)

6(W + 7) - 3 (W + 7) = 0
6W + 42 - 3W - 21 = 0
3W + 21 = 0
3W = -21
W = -7